# Rigorously Characterizing Dynamics with Machine Learning

### v1

@@ 2.1. Continuous Dynamics
We consider continuous functions $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and restrict our attention to the dynamics on a finite union of closed, bounded, convex $d$-dimensional polytopes $X\subset{\mathbb{R}}^{d}$ for which $f(X)\subset\mathop{\mathrm{int}}\nolimits(X)$. However, the discussion of this section applies to continuous maps on compact metric spaces [4, 30].
@@
The set of all attractors of $f$ is denoted by ${\mathsf{Att}}(f)$ and forms a bounded distributive lattice under the operations $\vee$=$\cup$ = ∪ and $A_{1}\wedge A_{2}=\omega(A_{1}\cap A_{2})$ [18] (for basic definitions associated with lattices see Section 7). The minimal element is ${\mathbf{0}}=\emptyset$ and the maximal element is ${\mathbf{1}}=\omega(X,f)$. Given $A\in{\mathsf{Att}}(f)$, the dual repeller to $A$ is defined to be $A^{*}:=\{x\in{\mathbf{1}}\mid\omega(x,f)\cap A=\emptyset\}$.
@@
${\mathbf{1}}_{\theta}:=\omega([-2,2],f_{\theta})=\begin{cases}\left\{{0}\right\}&\text{if $\theta<1$}\\ [0,\theta]&\text{if $1\leq\theta$.}\end{cases}$
@@
Conley’s decomposition theorem [4, 37] states that
(2) $CR(f)=\bigcap_{A\in{\mathsf{Att}}(f)}A\cup A^{*}$
is an invariant set, called the chain recurrent set. Furthermore, all recurrent dynamics of $f$ occurs in $CR(f)$ and $f$ exhibits gradient-like dynamics in the complement of $CR(f)$.
In general ${\mathsf{Att}}(f)$ may be countably infinite [4]. With this in mind, we restrict our attention to finite sublattices of ${\mathsf{Att}}(f)$ containing both ${\mathbf{0}}$ and ${\mathbf{1}}$. We leave it to the reader to check that if ${\mathsf{A}}$ is such a finite sublattice, then $CR(f)\subset\bigcap_{A\in{\mathsf{A}}}A\cup A^{*}$.
@@
where ${\mathsf{J}}^{\vee}({\mathsf{A}})$ denotes the join-irreducible elements of ${\mathsf{A}}$ and $A^{<}$ denotes the immediate predecessor of $A\in{\mathsf{J}}^{\vee}({\mathsf{A}})$ [20]. The elements of ${\mathsf{MR}}({\mathsf{A}})$, called Morse sets, are nonempty, pairwise disjoint invariant sets.
@@
As presented above, a finite sublattice of ${\mathsf{Att}}(f)$ gives rise to a well defined Morse representation. The converse is also true, i.e., by [20, Theorem 6] given a Morse representation ${\mathsf{MR}}$ of $f$ there is a unique finite lattice of attractors ${\mathsf{A}}$ such that (3) holds. In particular, (3) gives rise to a poset isomorphism defined by
@@
${\mathsf{MR}}({\mathsf{Att}}(f_{\theta}),\leq_{f_{\theta}})=\begin{cases}0&\text{if $\theta<1$}\\ (\left\{{0,1}\right\},0<_{f_{\theta}}1)&\text{if $\theta=1$}\\ (\left\{{0,1,\theta}\right\},0<_{f_{\theta}}1>_{\theta}\theta)&\text{if $\theta>1$.}\end{cases}$
@@
As Examples 2.6 demonstrates, Morse representations need not be stable with respect to perturbations, and thus, cannot be expected to be computable in general.
The partial order associated with a Morse representation provides a characterization of the structure of the non-recurrent dynamics. To identify finer structures we make use of the Conley index that we introduce via an extremely terse series of definitions. For more detailed discussion we refer the reader to [4, 36, 32, 41, 10, 30, 31].
@@
For the definition of shift equivalence, the reader is referred to [23, 2, 10, 31]. Over fields, the homology Conley index can be characterized in a more simple way: it is determined by the rational canonical form of $f_{P*}$, excluding nilpotent blocks [31]. See [23, 7.3–7.5] for an explanation of this characterization. Examples of computing the homology Conley index using rational canonical form are contained in [3] and use an efficient algorithm due to Storjohann [40]. See also [31] for a discussion of the limitations of using field coefficients. For the purposes of this paper, it is sufficient to know that in the context of the homology Conley index, shift equivalence is computable [31].
@@ 2.2. Computable Objects
Let ${\mathsf{ABlock}}(X,f)$ denote the set of all attracting blocks for $f$. The following proposition follows from the work of [18]. We provide an elementary proof to make this paper more self-contained.
@@
Let $X$ be a regular compact subset of ${\mathbb{R}}^{d}$. Let $f$ be a continuous function such that $f(X)\subset\mathop{\mathrm{int}}\nolimits(X)$. Then, the set ${\mathsf{ABlock}}(X,f)$ is a bounded, distributive lattice with respect to $\vee$=$\cup$ = ∪ and $\wedge$=$\cap$ = ∩. The neutral elements of ${\mathsf{ABlock}}(X,f)$ are ${\mathbf{1}}_{b}=X$ and ${\mathbf{0}}_{b}=\varnothing$.
@@
Let ${\mathsf{A}}$ be a finite sublattice of ${\mathsf{Att}}(f)$. By [19, Theorem 1.2] there exists a sublattice ${\mathsf{N}}$ of ${\mathsf{ABlock}}(f)$ and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$ for all $A\in{\mathsf{A}}$. Furthermore, consider a Morse representation ${\mathsf{MR}}({\mathsf{A}})$ and the corresponding isomorphism $\psi\colon{\mathsf{MR}}({\mathsf{A}})\to{\mathsf{J}}^{\vee}({\mathsf{A}})$ of Remark 2.5. Then, for every $M\in{\mathsf{MR}}({\mathsf{A}})$, $M=\mathop{\mathrm{Inv}}\nolimits(k(\psi(M))\setminus k(\psi(M)^{<}),f)$.
@@
is a sublattice of ${\mathsf{ABlock}}(f_{1.5})$ and $k\colon{\mathsf{Att}}(f_{1.25})\to{\mathsf{N}}$ generated by $k(0)=[-2,0.25]$ and $k(\theta)=[1.25,2]$ is a lattice isomorphism. Furthermore, $0=\mathop{\mathrm{Inv}}\nolimits([-2,0.25],f_{1.5})$, $\theta=\mathop{\mathrm{Inv}}\nolimits([1.25,2],f_{1.5})$, and $1=\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus([-2,0.25]\cup[1.25,2]),f_{1.5})$.
@@
This stability is exhibited in Example 2.15 where ${\mathsf{N}}$ is a lattice of ${\mathsf{ABlock}}(f_{\theta})$ for $1.751.25<\theta<$ 1.75.
The following result follows from [36, Corollary 4.4].
@@
${\mathsf{MTile}}({\mathsf{N}}):=\left\{{T(N):=\mathop{\mathrm{cl}}\nolimits(N\setminus N^{<})}\,\left|\,{N\in{\mathsf{J}}^{\vee}({\mathsf{N}})}\right.\right\}.$
@@ A Morse decomposition of $f$ is an order-embedding
Consider a finite sublattice ${\mathsf{N}}\subset{\mathsf{ABlock}}(f)$. Define $({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f}})$ by
${\mathsf{MR}}=\left\{{M(N):=\mathop{\mathrm{Inv}}\nolimits(N\setminus N^{<},f)\neq\emptyset}\,\left|\,{N\in{\mathsf{J}}^{\vee}({\mathsf{N}})}\right.\right\},$
@@
We use ${\mathsf{MTile}}({\mathsf{N}})$ to compute the Conley indices of $\mathop{\mathrm{Inv}}\nolimits([-2,0.25],f_{1.5})$, $\mathop{\mathrm{Inv}}\nolimits([1.25,2],f_{1.5})$, and $\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus([-2,0.25]\cup[1.25,2]),f_{1.5})$. In particular, using field coefficients ${\mathbb{F}}$
@@
Observe that unlike in Example 2.23, we cannot use the Conley index to identify that $[1,1.5]=\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus[-2,0.25],f_{1.5})$ is nonempty.
@@ 2.3. Summary of Theory
Let $({\mathsf{MR}}_{f},\leq_{{\mathsf{MR}}_{f}})$ be a Morse representation of $f$ with minimal partial order $\leq_{{\mathsf{MR}}_{f}}$. By [20, Theorem 6] there exists a unique finite sublattice of ${\mathsf{Att}}(f)$ and a poset isomorphism $\psi\colon({\mathsf{MR}}_{f},\leq_{{\mathsf{MR}}_{f}})\to({\mathsf{J}}^{\vee}({\mathsf{A}}),\subset)$. By [19, Theorem 1.2] there exists a sublattice ${\mathsf{N}}$ of ${\mathsf{ABlock}}(f)$ and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$. This implies there is a poset isomorphism $\kappa:{\mathsf{J}}^{\vee}({\mathsf{A}})\to{\mathsf{J}}^{\vee}({\mathsf{N}})$.
@@
As indicated by Remark 2.16, for each $N\in{\mathsf{N}}$ there exists $\epsilon_{N}>0$ such that if $\sup_{x\in X}\|f(x)-g(x)\|<\epsilon_{N}$, then $N$ is an attracting block for $g$. Set $\epsilon=\min\left\{{\epsilon_{N}}\,\left|\,{N\in{\mathsf{N}}}\right.\right\}$. Then ${\mathsf{N}}$ is a finite sublattice of ${\mathsf{ABlock}}(g)$. By Proposition 2.22 we have the existence of a Morse decomposition $\mu\colon{\mathsf{MR}}_{g}\to{\mathsf{J}}^{\vee}({\mathsf{N}})$ where
${\mathsf{MR}}_{g}=\left\{{M(N):=\mathop{\mathrm{Inv}}\nolimits(N\setminus N^{<},g)\neq\emptyset}\,\left|\,{N\in{\mathsf{J}}^{\vee}({\mathsf{N}})}\right.\right\}.$
@@
Let $M_{g}\in{\mathsf{MR}}_{g}$. Then by definition, $M_{g}=\mathop{\mathrm{Inv}}\nolimits(N\setminus N^{<},g)$ for some $N\in{\mathsf{J}}^{\vee}({\mathsf{N}})$. Since $\mathop{\mathrm{cl}}\nolimits(N\setminus N^{<})$ is an isolating neighborhood for all functions $g\colon X\to X$ such that $\sup_{x\in X}\|f(x)-g(x)\|<\epsilon$, by [10]
@@ 3. Combinatorial Dynamics
As shown in [19], ${\mathsf{Invset}}^{+}({\mathcal{F}})$ is a distributive lattice under the operations $\vee$=$\cup$ = ∪ and $\wedge$=$\cap$ = ∩, with minimal element ${\mathbf{0}}=\emptyset$ and maximal element ${\mathbf{1}}={\mathcal{X}}$.
@@
Define the relation $\sim$ on ${\mathcal{X}}$ by $\xi\sim\xi^{\prime}$ if there exists walks $\xi{\rightsquigarrow}\xi^{\prime}$ and $\xi^{\prime}{\rightsquigarrow}\xi$. We note that $\sim$ is an equivalence relation. A directed graph ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ is strongly connected if the quotient set ${\mathcal{X}}/\sim$ consists of one equivalence class, meaning there exists a walk between every pair of vertices.
Given an arbitrary directed graph ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$, the strongly connected components are the induced subgraphs with vertex sets equal to equivalence classes of ${\mathcal{X}}/\sim$. Equivalently, the strongly connected components are the maximal strongly connected subgraphs of ${\mathcal{F}}$. Various linear time algorithms for computing strongly connected components have been developed: for example, Tarjan’s algorithm [43] and Kosaraju-Sharir’s algorithm [1, 38].
[…]

### v2

@@ 2.1. Continuous Dynamics
We consider continuous functions $f\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and restrict our attention to the dynamics on a finite union of closed, bounded, convex $d$-dimensional polytopes $X\subset{\mathbb{R}}^{d}$ for which $f(X)\subset\mathop{\mathrm{int}}\nolimits(X)$. However, the discussion of this section applies to continuous maps on compact metric spaces [conley:cbms, mischaikow:mrozek].
@@
The set of all attractors of $f$ is denoted by ${\mathsf{Att}}(f)$ and forms a bounded distributive lattice under the operations $\vee=\cup$ and $A_{1}\wedge A_{2}=\omega(A_{1}\cap A_{2})$ [kalies:mischaikow:vandervorst:14] (for basic definitions associated with lattices see Section SM1 in the Supplementary Materials). The minimal element is $\mathbf{0}$$\emptyset$ and the maximal element is ${\mathbf{1}}=\omega(X,f)$. Given $A\in{\mathsf{Att}}(f)$, the dual repeller to $A$ is defined to be $A^{*}:=\{x\in{\mathbf{1}}\mid\omega(x,f)\cap A=\emptyset\}$.
@@
[…]
@@
Conley’s decomposition theorem [conley:cbms, robinson] states that $CR(f)=\bigcap_{A\in{\mathsf{Att}}(f)}A\cup A^{*}$ is an invariant set, called the chain recurrent set. Furthermore, all recurrent dynamics of $f$ occurs in $CR(f)$ and $f$ exhibits gradient-like dynamics in the complement of $CR(f)$.
In general ${\mathsf{Att}}(f)$ may be countably infinite [conley:cbms]. With this in mind, we restrict our attention to finite sublattices of ${\mathsf{Att}}(f)$ containing both $\mathbf{0}$} and $\mathbf{1}$}. We leave it to the reader to check that if $\mathsf{A}$} is such a finite sublattice, then $CR(f)\subset\bigcap_{A\in{\mathsf{A}}}A\cup A^{*}$.
@@
where ${\mathsf{J}}^{\vee}({\mathsf{A}})$ denotes the join-irreducible elements of $\mathsf{A}$} and $A^{<}$ denotes the immediate predecessor of $A\in{\mathsf{J}}^{\vee}({\mathsf{A}})$ [kalies:mischaikow:vandervorst:21]. The elements of ${\mathsf{MR}}({\mathsf{A}})$, called Morse sets, are nonempty, pairwise disjoint invariant sets.
@@
As presented above, a finite sublattice of ${\mathsf{Att}}(f)$ gives rise to a well defined Morse representation. The converse is also true, i.e., by [kalies:mischaikow:vandervorst:21, Theorem 6] given a Morse representation $\mathsf{MR}$} of $f$ there is a unique finite lattice of attractors $\mathsf{A}$} such that (2) holds. In particular, (2) gives rise to a poset isomorphism defined by
@@
${\mathsf{MR}}({\mathsf{Att}}(f_{\theta}),\leq_{f_{\theta}})=\begin{cases}0&\text{if $\theta<1$}\\ (\left\{{0,1}\right\},0<_{f_{\theta}}1)&\text{if $\theta=1$}\\ (\left\{{0,1,\theta}\right\},0<_{f_{\theta}}1>_{f_{\theta}}\theta)&\text{if $\theta>1$.}\end{cases}$
@@
As Example 2.6 demonstrates, Morse representations need not be stable with respect to perturbations, and thus, cannot be expected to be computable in general.
The partial order associated with a Morse representation provides a characterization of the structure of the non-recurrent dynamics. To identify finer structures we make use of the Conley index that we introduce via an extremely terse series of definitions. For more detailed discussion we refer the reader to [conley:cbms, robbin:salamon, mrozek:99, szymczak:95, franks:richeson, mischaikow:mrozek, mischaikow:weibel].
@@
For the definition of shift equivalence, the reader is referred to [lind:marcus, boyle, franks:richeson, mischaikow:weibel]. Over fields, the homology Conley index can be characterized more simply by the rational canonical form of $f_{P*}$, excluding nilpotent blocks [mischaikow:weibel], [lind:marcus, 7.3–7.5]. Examples of computing the homology Conley index are contained in [bush:cowan:harker:mischaikow] and use an efficient algorithm due to Storjohann [storjohann]. For the purposes of this paper, it is sufficient to know that in the context of the homology Conley index, shift equivalence is computable [mischaikow:weibel].
@@ 2.2. Computable Objects
Let ${\mathsf{ABlock}}(X,f)$ denote the set of all attracting blocks for $f$. The following proposition follows from the work of [kalies:mischaikow:vandervorst:14]. We provide an elementary proof to make this paper more self-contained.
@@
Let $X$ be a regular compact subset of ${\mathbb{R}}^{d}$. Let $f$ be a continuous function such that $f(X)\subset\mathop{\mathrm{int}}\nolimits(X)$. Then, the set ${\mathsf{ABlock}}(X,f)$ is a bounded, distributive lattice with respect to $\vee=\cup$ and $\wedge=\cap$. The neutral elements of ${\mathsf{ABlock}}(X,f)$ are ${\mathbf{1}}_{b}=X$ and ${\mathbf{0}}_{b}=\varnothing$.
@@
Let $\mathsf{A}$} be a finite sublattice of ${\mathsf{Att}}(f)$. By [kalies:mischaikow:vandervorst:15, Theorem 1.2] there exists a sublattice $\mathsf{N}$} of ${\mathsf{ABlock}}(f)$ and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$ for all $A\in{\mathsf{A}}$. Furthermore, consider a Morse representation ${\mathsf{MR}}({\mathsf{A}})$ and the corresponding isomorphism $\psi\colon{\mathsf{MR}}({\mathsf{A}})\to{\mathsf{J}}^{\vee}({\mathsf{A}})$ of Remark 2.5. Then, for every $M\in{\mathsf{MR}}({\mathsf{A}})$, $M=\mathop{\mathrm{Inv}}\nolimits(k(\psi(M))\setminus k(\psi(M)^{<}),f)$.
@@
is a sublattice of ${\mathsf{ABlock}}(f_{1.5})$ and $k\colon{\mathsf{Att}}(f_{1.25})\to{\mathsf{N}}$ generated by k(0)=[2,0.25]k(0)=[-2,0.25] and $k(\theta)=[1.25,2]$ is a lattice isomorphism. Furthermore, $0=\mathop{\mathrm{Inv}}\nolimits([-2,0.25],f_{1.5})$, $\theta=\mathop{\mathrm{Inv}}\nolimits([1.25,2],f_{1.5})$, and $1=\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus([-2,0.25]\cup[1.25,2]),f_{1.5})$.
@@
This stability is exhibited in Example 2.15 where $\mathsf{N}$} is a lattice of ${\mathsf{ABlock}}(f_{\theta})$ for $1.25<\theta<1.75$.
The following result follows from [robbin:salamon, Corollary 4.4].
@@
[…]
@@ A Morse decomposition of $f$ is an order-embedding
Consider a finite sublattice ${\mathsf{N}}\subset{\mathsf{ABlock}}(f)$. Define
$({\mathsf{MR}},\leq_{{\mathsf{MR}}_{f}})$ by
[…]
@@
We use ${\mathsf{MTile}}({\mathsf{N}})$ to compute the Conley indices of $\mathop{\mathrm{Inv}}\nolimits([-2,0.25],f_{1.5})$,
$\mathop{\mathrm{Inv}}\nolimits([1.25,2],f_{1.5})$, and $\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus([-2,0.25]\cup[1.25,2]),f_{1.5})$. In particular, using field coefficients $\mathbb{F}$}
@@
Observe that unlike in Example 2.23, we cannot use the Conley index to identify that $1,1.5]=\mathop{\mathrm{Inv}}\nolimits([-2,2]\setminus[-2,0.25],f_{1.5})$ is nonempty.
@@ 2.3. Summary of Theory
Let $({\mathsf{MR}}_{f},\leq_{{\mathsf{MR}}_{f}})$ be a Morse representation of $f$ with minimal partial order $\leq_{{\mathsf{MR}}_{f}}$. By [kalies:mischaikow:vandervorst:21, Theorem 6] there exists a unique finite sublattice of ${\mathsf{Att}}(f)$ and a poset isomorphism $\psi\colon({\mathsf{MR}}_{f},\leq_{{\mathsf{MR}}_{f}})\to({\mathsf{J}}^{\vee}({\mathsf{A}}),\subset)$. By [kalies:mischaikow:vandervorst:15, Theorem 1.2] there exists a sublattice $\mathsf{N}$} of ${\mathsf{ABlock}}(f)$ and a lattice isomorphism $k\colon{\mathsf{A}}\to{\mathsf{N}}$ such that $\omega(k(A),f)=A$. This implies there is a poset isomorphism $\kappa:{\mathsf{J}}^{\vee}({\mathsf{A}})\to{\mathsf{J}}^{\vee}({\mathsf{N}})$.
@@
[…]
@@
Let $M_{g}\in{\mathsf{MR}}_{g}$. Then by definition, $M_{g}=\mathop{\mathrm{Inv}}\nolimits(N\setminus N^{<},g)$ for some $N\in{\mathsf{J}}^{\vee}({\mathsf{N}})$. Since $\mathop{\mathrm{cl}}\nolimits(N\setminus N^{<})$ is an isolating neighborhood for all functions $g\colon X\to X$ such that $\sup_{x\in X}\|f(x)-g(x)\|<\epsilon$, by [franks:richeson]
@@ 3. Combinatorial Dynamics
As shown in [kalies:mischaikow:vandervorst:15], ${\mathsf{Invset}}^{+}({\mathcal{F}})$ is a distributive lattice under the operations $\vee=\cup$ and $\wedge=\cap$, with minimal element $\mathbf{0}$$\emptyset$ and maximal element ${\mathbf{1}}={\mathcal{X}}$.
@@
Define the relation $\sim$ on $\mathcal{X}$} by $\xi\sim\xi^{\prime}$ if there exists walks $\xi{\rightsquigarrow}\xi^{\prime}$ and $\xi^{\prime}{\rightsquigarrow}\xi$. We note that $\sim$ is an equivalence relation. A directed graph ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ is strongly connected if the quotient set $\mathcal{X}$$\sim$ consists of one equivalence class, meaning there exists a walk between every pair of vertices.
Given an arbitrary directed graph ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$, the strongly connected components are the induced subgraphs with vertex sets equal to equivalence classes of $\mathcal{X}$$\sim$. Equivalently, the strongly connected components are the maximal strongly connected subgraphs of $\mathcal{F}$}. Various linear time algorithms for computing strongly connected components have been developed: for example, Tarjan’s algorithm [tarjan] and Kosaraju-Sharir’s algorithm [aho:hopcroft:ullman, sharir].
Let $\pi\colon{\mathcal{X}}\to{\mathsf{SCC}}({\mathcal{F}}):={\mathcal{X}}/\sim$ denote the quotient map from $\mathcal{X}$} to the equivalence classes of strongly connected components. The graph obtained by contracting edges that have incident vertices in the same strongly connected component is the condensation graph of ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$, denoted by $\bar{{\mathcal{F}}}:{\mathsf{SCC}}({\mathcal{F}})\rightrightarrows{\mathsf{SCC}}({\mathcal{F}})$. The condensation graph is an acyclic directed graph [corman:leiserson:rivest:stein], and it may be viewed as a compression of $\mathcal{F}$} that retains reachability information. Due to acyclicity, ${\mathsf{SCC}}({\mathcal{F}})$ is a poset under the relation $\zeta^{\prime}\leq_{{\mathsf{SCC}}({\mathcal{F}})}\zeta$ if $\zeta{\rightsquigarrow}\zeta^{\prime}$ in $\bar{{\mathcal{F}}}$. A recurrent component of ${\mathcal{F}}:{\mathcal{X}}\rightrightarrows{\mathcal{X}}$ is a strongly connected component with at least one edge in $\mathcal{F}$}.